OpenStudy (anonymous):

If f(x) = x2 - x + 2, then find the value for f(-3). can you please just solve the whole problem and show your work this is worth ten percent of my grade and if i fail i wil fail math i will fan you and give medals and mark you as best answer please :(

OpenStudy (johnweldon1993):

We cannot just give you the answer....but here is a quick hint f(x) f(-3) what's different between those two? well...-3...is in the place of x so in terms of your equations where you have f(x)= x² - x + 2 f(3) = (-3)² - (-3) + 2 you literally just plug -3 in for every 'x' that you see...can you solve this?

OpenStudy (whpalmer4):

Okay, here's the equation: \[f(x) = x^2 - x + 2\] You need to evaluate \(f(-3)\). To do that, you rewrite the equation substituting (-3) everywhere you see \(x\).

OpenStudy (whpalmer4):

Remember, a negative number * a negative number gives you a positive number.

OpenStudy (anonymous):

\[f(x) =x^2-x+2\]

OpenStudy (anonymous):

f(-3)=(-3)^2 -(-3)+2 f(-3)=9+3+2 f(-3)=14

OpenStudy (anonymous):

this is the answer.. write in order that I wrote..

OpenStudy (anonymous):

f(-3)=(-3)^2 -(-3)+2 f(-3)=9+3+2 f(-3)=14 so 14 is the answer?

OpenStudy (anonymous):

yes that is the exact answer.. trust :)

OpenStudy (whpalmer4):

Yes, I agree that f(-3) = 14.

OpenStudy (whpalmer4):

John showed you exactly what you needed to do, 40 minutes ago...

OpenStudy (anonymous):

thanks that ewas just alot more simple lol:)

OpenStudy (whpalmer4):

Yeah, simpler because he did the whole thing for you, violating the OpenStudy code of conduct...